How long will it take an investment of $2000 to double if the investment earns interest at the rate of 5%/year compounded monthly?
Answer by mananth(16946)

Find number of years
Principal P = 2000
Amount= 4000
years=n
compounded (t) 12 (monthly)
Rate = 5 (5%/12) -> 0.004166667
Amount = P*((n+r)/n)^n*t

4000 = 2000 *( 1 + 0.004166667 )^ n*t
2 = *( 1 + 0.004166667 )^ n* 12
ln 2 = 12 n *ln 1.004166667
0.693147181 = 12 0.00415801 *n
13.89180473 = n
Number of years = 13.89